# Enhancing Zero-Shot Facial Expression Recognition by LLM Knowledge Transfer

### v2

@@ IV-C Comparisons with the State-of-the-Art
TABLE III: Zero-shot FER results with different prompts in comparison with CLIP model. Both single prompts and prompt ensembles are investigated. We employ three types for single prompts, labelled as P-1, P-2, and P-3 respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations, labelled as P-4 and P-5.
@@
BLIP2 [23] is a state-of-the-art model on various vision-language tasks, to test its zero-shot performance on the FER task, inspired by work [22], we take the facial images as input and prompt BLIP2 by using: “Question: Play the role of a facial expression classification expert and select one category with the highest similarity to the input image. Here are the optional categories: ‘contempt’, ‘fear’, ‘disgust’, ‘surprise’, ‘anger’, ‘neutral’, ‘sadness’, ‘happiness’. Answer:”. We then calculate the WAR and UAR based on the output. As shown in Tab. I, our results outperform the BLIP2 model on all the FER test sets. Specifically, our ViT-B-16-based zero-shot results can beat BLIP2 in most SFER test sets.
Furthermore, we also compare our results with several other visual-language models. VideoCLIP [43], XCLIP [44], and EmotionCLIP [27] are proposed for video analysis, video recognition, and video emotion recognition, respectively, in which VideoCLIP was pre-trained on video clips and text transcriptions, XCLIP was pre-trained on action recognition dataset with action labels, and EmotionCLIP was pre-trained on the videos and conversations. These models all learn spatial-temporal features during the training. However, our method with the bare static model can still outperform them on four DFER test sets. Moreover, we also compared our method with CLIP betterment: FaRL and FLAVA, in which the FaRL model was pre-trained on LAION-FACE specializing in the face analysis task and FLAVA improves CLIP zero-shot performance by introducing both unimodal and multimodal architecture. The results in Tab. I demonstrate our results are superior to these two methods on all of the FER test sets. Compared with the EmoCLIP [28], which was trained on the MAFW clips and facial action descriptions, our method outperforms it on DFEW, FERV39k, and AFEW test sets in five metrics without explicit text descriptions, indicating the effectiveness of the proposed method.
IV-D Ablation Analysis
Evaluation of the Projection Head Effectiveness: Our method is built on the CLIP model [20], to evaluate the effectiveness of the learned projection head, we compare our method to the CLIP model from two views: various ViT models and various zero-shot prompts. Our method uses the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”. As shown in Tab. II, under different ViT models, our method outperforms the plain CLIP model on all the test sets. With ViT-B-32, our method improves the average UAR by 2.71% and average WAR by 4.42% over eight test sets. With ViT-B-16, our method improves the average UAR by 5.47% and average WAR by 6.75% over eight test sets. With ViT-L-14, our method improves the average UAR by 8.07% and average WAR by 8.23% over eight test sets.
Furthermore, we compare our method with CLIP using different prompts, in which both single prompt and prompt ensembles are considered. Specifically, for a single prompt, we utilize “{class name}.”, “an expression of {class name}.”, and “a photo of a face with an expression of {class name}.”, respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations. All the prompt templates are outlined below:
TABLE IV: Zero-shot FER results of our method with different instructions. Both task-related and task-unrelated instructions are investigated. All models use the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.
@@ IV-D Ablation Analysis
For each prompt, we compare our method with the CLIP model one by one. As shown in Tab. III, except for seven values slightly lower than the CLIP model, our method outperforms the CLIP model on all the rest FER test sets, demonstrating the effectiveness of the proposed method. It should be noted that, compared with the CLIP model, only one extra projection matrix is added on top of the two encoders in our method and this projection matrix was trained in an unsupervised manner.
Evaluation of Different Instructions: To verify the effectiveness of the instruction adopted in our method, we conducted experiments on two types of instructions: task-related and task-unrelated. Regarding the task-unrelated instructions, we consider both empty and random text as input and take the average of them as the final results. For the task-related instruction, we provide results with three instructions and their average, including “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”, “Please play the role of a facial expression recognition expert. Describe the facial expression of the person in the image.”, and “Please describe the detailed facial actions of the person in the image.”.
As shown in Tab. IV, the task-related instructions are superior to task-unrelated instructions on both SFER and DFER, in which there are 3.14% average UAR improvements and 4.02% average WAR improvements over eight test sets with ViT-B-32, 2.87% average UAR improvements and 4.86% average WAR improvements over eight test sets with ViT-B-16, and 1.38% average UAR improvements and 1.87% average WAR improvements over eight test sets with ViT-L-14. In addition, the three task-related instructions show a small variance in downstream zero-shot prediction. We believe that task-related instructions facilitate LLMs to generate semantic features with task-aware ability, providing better objects for projection head optimisation.
TABLE V: Zero-shot FER results of our method with different projection heads. All models are based on ViT-L-14, and use the instruction “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.” and the prompt “a photo of a face with an expression of [class].” for zero-shot prediction. 2-MLP and 3-MLP denote 2-layer and 3-layer multilayer perceptron, respectively. TE denotes Transformer encoder. LM denotes our learnable linear matrix.
Projection Heads Static FER (UAR/WAR) Dynamic FER (UAR/WAR)
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
TABLE VI: Zero-shot FER results of varied image and text encoders. PH denotes the projection head trained in our method.
Image Encoder Text Encoder Static FER Dynamic FER
RAF-DB AffectNet-7 AffectNet-8 FERPlus DFEW FERV39k MAFW AFEW
Figure 3: Comparison of the video-to-text retrieval performance on the subset of the MAFW.
Figure 5: Normalized CLIP latent visual feature variance of each emotion on FERPlus.
Evaluation of Different Projection Heads: In our method, to align the original CLIP feature representations into task-specific LLM knowledge space, we propose to utilize a projection head as a linear mapping to fulfil this. To investigate if the improvement of the complexity of the projectors can further improve the alignment, we also conduct experiments with 2-layer MLP, 3-layer MLP, and Transformer Encoders.
As shown in Tab. V, even with a simply learnable matrix and fewer parameters, our strategy is more powerful and efficient. Compared to the Transformer Encoders, except for the WAR results on FERV39k, ours shows much better results than more complex Transformer Encoders. We believe that more complex non-linear models will entail overfitting to the LLMs feature space, performing worse in downstream zero-shot FER tasks. The linear projection head adopted in our method is more suitable and interpretable for mapping the original CLIP feature space to a task-specific one, maintaining the capacity of the existing model to capture robust, versatile representations while tailoring it to the FER domain with minimal trainable parameters.
Evaluation of Vision-to-Text-Description Understanding Ability: To verify the vision-to-text-description ability of our method, we conducted experiments using facial videos and their corresponding facial action descriptions. Specifically, we performed a video-text retrieval task on a subset of the MAFW dataset [49]. This subset comprises 8,034 sample-level text descriptions, each detailing specific facial actions. We utilized Precision@k as the metric for video-to-text retrieval and compared the performance of our method against several other vision-language models. As illustrated in Fig. 3, our Exp-CLIP model outperforms the other methods, which means our model retrieves more relevant text descriptions for a given facial video compared to other models and excels in understanding and retrieving complex and detailed facial action descriptions.
Evaluation of Varied Image and Text Encoders: In our method, we proposed to project the original CLIP feature space into a task-aware space, in which the LLM is adopted as a guide. Essentially, on top of the CLIP model, the proposed method is trying to distillate task-specific knowledge from the LLM encoder. To investigate if the LLM can do better than our method on zero-shot FER tasks, we conducted experiments using the LLM as either an image encoder (with I2T module), text encoder, or both. As shown in Tab. VI, adopting the LLM as either an image encoder or text encoder results in a worse zero-shot preference. Even when using the LLM as both an image and text encoder, the results remain undesirable and significantly lower than the plain CLIP model and our method. We believe the LLM excels in the semantic perception of images but performs poorly in recognition tasks. By contrast, based on the CLIP model, our method is more efficient and accurate in zero-shot FER task.
IV-E Visualizations
To prove the benefit of the enhanced CLIP visual features, we utilize t-SNE [53] to illustrate the feature distribution of the original CLIP feature representations and projected feature representations. Different from the supervised methods learned with annotations which can gather samples of the same category compactly, zero-shot visual feature representations are learned by aligning the cross-modality representations, presenting a weak performance in gathering samples compactly. Despite this, from the visual feature distributions in Fig. 4, we can see the samples of the same category can be better gathered after the projection, especially for emotions of happiness and disgust. This demonstrates the model learns more discriminative latent visual features after the projection head, and can also explain why our method obtains superior results to the original CLIP model. In addition, we also calculate the normalized feature variance of each emotion on FERPlus, as shown in Fig. 5, the projected features demonstrate a lower variance than the plain CLIP feature space, proving the projected features are more concentrated.
((a))
((b))
((c))
((d))
((e))
((f))
@@ IV-E Visualizations
Figure 6: Confusion matrix compared with CLIP model on static FER datasets.
((a))
((b))
((c))
((d))
((e))
((f))
@@
Figure 7: Confusion matrix compared with CLIP model on dynamic FER datasets.
@@
Moreover, to offer a more extensive comparison with the CLIP model, we include the learned confusion matrices of both our method and the CLIP model, displayed in Fig. 6 and Fig. 7. From the confusion matrix on both static FER and dynamic FER datasets, we can see that our method learned better representations for each expression category, resulting in a more balanced accuracy across all categories. Specifically, our method can better distinguish between neutral expressions and other expressions. This suggests that our method captures more nuanced facial expression features derived from LLMs.
V Conclusion
Acknowledgments
This research utilised Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT. We acknowledge the assistance of the ITS Research team at Queen Mary University of London. Zengqun Zhao is funded by Queen Mary Principal’s PhD Studentships. We thank Dr. Ziquan Liu, James Oldfield, and Yimeng Gu for their valuable comments.

### v3

@@ A Comparison with CLIP on Varied Prompts
Table A: Zero-shot FER results with different prompts in comparison with CLIP model. Both single prompts and prompt ensembles are investigated. We employ three types for single prompts, labelled as P-1, P-2, and P-3 respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations, labelled as P-4 and P-5.
@@
Table B: Zero-shot FER results of our method with different instructions. Both task-related and task-unrelated instructions are investigated. All models use the prompt “a photo of a face with an expression of [class].” for zero-shot prediction.
@@
We compare our method with CLIP using different prompts, in which both single prompt and prompt ensembles are considered. Specifically, for a single prompt, we utilize “{class name}.”, “an expression of {class name}.”, and “a photo of a face with an expression of {class name}.”, respectively. For prompt ensembles, we utilize 5-prompt and 10-prompt configurations. All the prompt templates are outlined below:
@@
For each prompt, we compare our method with the CLIP model one by one. As shown in Tab. A, except for seven values slightly lower than the CLIP model, our method outperforms the CLIP model on all the rest FER test sets, demonstrating the effectiveness of the proposed method. It should be noted that, compared with the CLIP model, only one extra projection matrix is added on top of the two encoders in our method and this projection matrix was trained in an unsupervised manner.
@@
Figure A: Confusion matrix compared with CLIP model on static FER datasets.
B Evaluation of Different Instructions
To verify the effectiveness of the instruction adopted in our method, we conducted experiments on two types of instructions: task-related and task-unrelated. Regarding the task-unrelated instructions, we consider both empty and random text as input and take the average of them as the final results. For the task-related instruction, we provide results with three instructions and their average, including “Please play the role of a facial action describer. Objectively describe the detailed facial actions of the person in the image.”, “Please play the role of a facial expression recognition expert. Describe the facial expression of the person in the image.”, and “Please describe the detailed facial actions of the person in the image.”.
As shown in Tab. B, the task-related instructions are superior to task-unrelated instructions on both SFER and DFER, in which there are 3.14% average UAR improvements and 4.02% average WAR improvements over eight test sets with ViT-B-32, 2.87% average UAR improvements and 4.86% average WAR improvements over eight test sets with ViT-B-16, and 1.38% average UAR improvements and 1.87% average WAR improvements over eight test sets with ViT-L-14. In addition, the three task-related instructions show a small variance in downstream zero-shot prediction. We believe that task-related instructions facilitate LLMs to generate semantic features with task-aware ability, providing better objects for projection head optimisation.
@@ B Evaluation of Different Instructions
Figure B: Confusion matrix compared with CLIP model on dynamic FER datasets.
C Confusion Matrix Comparison with CLIP
@@
To offer a more extensive comparison with the CLIP model, we include the learned confusion matrices of both our method and the CLIP model, displayed in Fig. A and Fig. B. From the confusion matrix on both static FER and dynamic FER datasets, we can see that our method learned better representations for each expression category, resulting in a more balanced accuracy across all categories. Specifically, our method can better distinguish between neutral expressions and other expressions. This suggests that our method captures more nuanced facial expression features derived from LLMs.
Figure C: Normalized CLIP latent visual feature variance of each emotion on FERPlus.
Figure D: Comparison of the video-to-text retrieval performance on the subset of the MAFW.
D Normalized Feature Variance
The normalized feature variance of each emotion on FERPlus is shown in Fig. C. The projected features demonstrate a lower variance than the CLIP feature space, proving the projected features are more concentrated.
E Evaluation of Vision-to-Text-Description Understanding Ability
To verify the vision-to-text-description ability of our method, we conducted experiments using facial videos and their corresponding facial action descriptions. Specifically, we performed a video-text retrieval task on a subset of the MAFW dataset. This subset comprises 8,034 sample-level text descriptions, each detailing specific facial actions. We utilized Precision@k as the metric for video-to-text retrieval and compared the performance of our method against several other vision-language models. As illustrated in Fig. D, our Exp-CLIP model outperforms the other methods, which means our model retrieves more relevant text descriptions for a given facial video compared to other models and excels in understanding and retrieving complex and detailed facial action descriptions.
F More Sample-Level Predictions and Analysis
We provide more sample-level predictions in the Fig. E. From the probability predictions on top images, we observe that our Exp-CLIP model correctly recognizes facial expressions with high confidence. However, our zero-shot method still makes some incorrect predictions, particularly when distinguishing between similar emotions like Anger and Disgust, as well as Surprise and Fear, especially on hard or ambiguous samples. These misclassifications likely arise from the subtle differences in facial expressions between these emotions, which can be difficult even for supervised models. Further refinement of our model or additional fine-tuning on few-shot labelled data may help mitigate these specific shortcomings.
Figure E: Zero-shot facial expression predictions of the proposed Exp-CLIP.